# Epistemic Power, Objectivity and Gender in AI Ethics Labor: Legitimizing Located Complaints

### v2

@@ Abstract.
By entrenching the epistemic power of quantification, dominant AI ethics practices—Model Cards and similar interventions—risk legitimizing AI ethics as a project in equal and opposite measure to which they delegitimize and marginalize embodied and lived experiences as legitimate parts of the same project. In response, I propose humble technical practices: quantified or technical practices which specifically seek to make their epistemic limits clear in order to flatten hierarchies of epistemic power.
journalyear: 2024copyright: rightsretainedconference: 2024 ACM Conference on Fairness, Accountability, and Transparency; June 3–6, 2024; Rio de Janeiro, Brazilbooktitle: 2024 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’24), June 3–6, 2023, Rio de Janeiro, Brazildoi: XXXXisbn: YYYYYY
@@ 1. Introduction
Integrating empirical interviews together with theory, this paper shows how the work of AI ethics is often understood as a “subjective” endeavor, and thus within the technology cultures in which this work sits, is cast as lower status when compared with the engineering work, seen as objective or rational. [^1] The definitions of objectivity and subjectivity have taken nearly polar opposite meanings across history (Daston and Galison, 2007, p. 31). The more familiar “new” senses of these words define so-called objective phenomena as not interpreted, and are thus taken to be “external to or independent of the mind” (Dictionary, 2023b), and subjective phenomena are “relating to the thinking subject … proceeding from or taking place within the individual consciousness or perception; having its source in the mind” (Dictionary, 2023a). In recognizing that all knowledge is situated and embodied, and thus not independent of the mind (Haraway, 1988), I reject this dichotomy even while recognizing it to be a social reality for many, including for many of my participants. This is predicted by Forsythe’s argument about the “deletion of the social”(Forsythe, 2001, p. 28), which she positioned as a detriment foremost to AI functionality, but in our context also operates to delete social knowledges from AI ethics discussions, and position such discussion as lower status. For women and other minoritized groups doing ethics work and disproportionately filling ethics roles, this interacts with documented ways in which their knowledge and work is devalued and delegitimized (West et al., 2019). In response, I show how some participants seek to legitimize ethics work by casting it as objective, quantitative and therefore authoritative. However, I show how other participants attempt to speak about ethics from their own situated, embodied perspective, but find these attempts delegitimized.
@@
The empirical basis for this paper is 75 interviews with technologists conducted remotely between early 2021 and late 2023, in a multi-sited approach (Williams et al., 2014). Purposive and snowball recruitment (Palinkas et al., 2015) helped reach participants in a variety of contexts: the majority work in companies ranging from startups to multinationals, many work on open source projects, several are academics, and a smaller group are activists seeking to question local development of carceral technology. Many inhabited and moved between multiple such categories. Participants were in four continents: the majority based in North America and Europe but also including those in Asia and Africa. The majority were doing AI work at some point in its “supply chain” (Cobbe et al., 2023; Widder and Nafus, 2023): spanning dataset curation, academic AI research, building new models, to deploying them in consumer or B2B contexts. Corporate participants were often required to engage with corporate AI ethics practices. This includes some who have accepted AI ethics as part of their work, whether through intrinsic motivation or as part of their assigned duties (e. g., Metcalf et al., 2019), and many who have not. Interview questions were tailored to context and did not adopt consistent predetermined structure(Seidman, [n. d.]): some portion of participants spoke to me about their own ethical concerns, often after raising them to their employers and seeing them go unaddressed. This study received IRB or other ethics approval as appropriate.
@@
The variety of contexts in which I investigate AI ethics practices allows us to understand how a variety of epistemological stances interact with the legitimacy of AI ethics claims, and what is rendered possible or not as a result of these stances. Participants were invited to self-identify their gender: 57 identified as men, 13 identified as women, four identified as non-binary, and one declined to answer. Most interviews lasted one hour (though ranging between 25 minutes and 3 hours), and were transcribed and analyzed using inductive thematic analysis (Braun and Clarke, 2006, 2019) [^3] In an inductive thematic analysis, themes are grounded in and emerge from the data, as opposed to deductive analysis where themes are decided a priori, and is properly a reflexive process, where the researcher reflects on their assumptions, training, values, and how this may shape analysis. See also: (Braun and Clarke, 2021, 2006)., under an interpretivist epistemological paradigm (Lincoln and Guba, 2000).444 Interpretivism, similar to its younger cousin constructivism (Schwandt, [n. d.]), understands that individuals are experts in interpreting their own experience, asks researchers to deeply understand these interpretations, and with both sharing the task of constructing wider meaning from these. To positivists, this is an unacceptable limitation, as no view from nowhere objectivity is possible. See also (Lincoln and Guba, 2000; Daston and Galison, 2007; Gebru, 2021; Haraway, 1988). My analysis is therefore inflected by my own positionality: I am a male-identified scholar in the United States with formal training in computer science, mentored by feminist STS scholars, with experience organizing against pervasive surveillance (Guo and Ryan-Mosley, 2023; Gulland and Tan, 2023) and carceral technology (Davidson, 2020; Kubit, 2022). It is from attempts to reconcile differences in language and legitimacy between the first and last of these–computer science and carceral tech activism–that I am motivated to write this paper, but my positionality necessarily limits those I can talk with and inflects the interpretations I arrive at, and others may arrive differently. In some ways, these locations represent the “stepping in and stepping out” characteristic of fieldwork (Powdermaker, 1966) in that some of my nonbinary and woman-identified participants construct their AI ethics practice as self-consciously feminist or even necessarily constructed in femme-only spaces, yet also, the predominantly technical and male backdrop in which AI ethics work sits paints me as part of the ingroup relative to the bulk of participants in this work: men in technical roles.
@@
2. The lower status of “ethics” work within AI cultures
This section demonstrates how my participants understand ethics work to be lower status relative to “real” AI work, but similar phenomena have been examined before. Past research shows how certain work in technology communities, like writing documentation or building community, is often done by women, and is less valued than writing code (Nafus, 2012). More recent work within AI has even shown differences between high status work on models versus lower status work on the dataset models are built from (Sambasivan et al., 2021), and recent manifestos have attacked “tech ethics” and related on the basis that it hinders “progress” on endeavors such as “Becoming Technological Supermen” (Andreessen, 2023). Forsythe shows how these status divides are inscribed in the language used by the AI practitioners she studied: they restricted their use of the word “‘work’ to what they think of as ‘real AI’: building systems”  (Forsythe, 2001, p. 26), set in contrast to the other kinds of work such as grant writing, meetings, and also the “knowledge elicitation” interviews seen as boring but necessary to “capture” knowledge to encode into systems.
@@ 6. Conclusions: towards humble technical practices
Epistemic power is bound up with, and often flowing from, other kinds including economic and political power. For example, Forsythe writes how the (bad) appropriation of ethnographic methods in positivist ways by untrained AI researchers had economic outcomes: “my own research funds have dried up” (Forsythe, 2001, p. 143). It is therefore those with other kinds of power, along lines of race, gender, or wealth, that should first share or cede epistemic power. Notably, people harmed by AI are often not paid nor recognized for sharing their lived experience of this harm in the ways that academics may be paid to collect it (Sloane et al., 2022). Therefore, I call for more ways to legitimize the sharing of lived experience in and on its own terms as valid AI ethics practice, and indeed, the preeminent and most true form of such practice. This must be listened to, and AI Ethics must be seen as a project to legitimize and amplify these on their own terms.

### v4

# Epistemic Power in AI Ethics Labor: Legitimizing Located Complaints

@@ Abstract.
By entrenching the epistemic power of quantification, dominant AI ethics practices—employing Model Cards and similar interventions—risk legitimizing AI ethics as a project in equal and opposite measure to which they marginalize lived experience as a legitimate part of the same project. In response, I propose humble technical practices: quantified or technical practices which specifically seek to make their epistemic limits clear in order to flatten hierarchies of epistemic power.
journalyear: 2024copyright: acmlicensedconference: The 2024 ACM Conference on Fairness, Accountability, and Transparency; June 3–6, 2024; Rio de Janeiro, Brazilbooktitle: The 2024 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’24), June 3–6, 2024, Rio de Janeiro, Brazildoi: 10.1145/3630106.3658973isbn: 979-8-4007-0450-5/24/06
@@ 1. Introduction
Integrating empirical interviews together with theory, this paper shows how the work of AI ethics is often understood as a “subjective” endeavor, and thus within the technology cultures in which this work sits, is cast as lower status when compared with the engineering work, seen as objective or rational. [^1] The definitions of objectivity and subjectivity have taken nearly polar opposite meanings across history (Daston and Galison, 2007, p. 31). The more familiar “new” senses of these words define so-called objective phenomena as not interpreted, and are thus taken to be “external to or independent of the mind” (Dictionary, 2023b), and subjective phenomena are “relating to the thinking subject … proceeding from or taking place within the individual consciousness or perception; having its source in the mind” (Dictionary, 2023a). In recognizing that all knowledge is situated and embodied, and thus not independent of the mind (Haraway, 1988), I reject this dichotomy even while recognizing it to be a social reality for many, including for many of my participants. This is predicted by Forsythe’s argument about the “deletion of the social” (Forsythe, 2001, p. 28), which she positioned as a detriment foremost to AI functionality, but in our context also operates to delete social knowledges from AI ethics discussions, and position such discussion as lower status. For women and other minoritized groups doing ethics work and disproportionately filling ethics roles, this interacts with documented ways in which their knowledge and work is devalued and delegitimized (West et al., 2019). In response, I show how some participants seek to legitimize ethics work by casting it as objective, quantitative and therefore authoritative. However, I show how other participants attempt to speak about ethics from their own situated, embodied perspective, but find their attempts delegitimized.
@@
The empirical basis for this paper is 75 interviews with technologists conducted remotely between early 2021 and late 2023, in a multi-sited approach (Williams et al., 2014). Purposive and snowball recruitment (Palinkas et al., 2015) helped reach participants in a variety of contexts: the majority work in companies ranging from startups to multinationals, many work on open source projects, several are academics, and a smaller group are activists seeking to question local development of carceral technology. Many inhabited and moved between multiple such categories. Participants were in four continents: the majority based in North America and Europe but also including those in Asia and Africa. The majority were doing AI work at some point in its “supply chain” (Cobbe et al., 2023; Widder and Nafus, 2023): spanning dataset curation, academic AI research, building new models, to deploying them in consumer or B2B contexts. Corporate participants were often required to engage with corporate AI ethics practices. This includes some who have accepted AI ethics as part of their work, whether through intrinsic motivation or as part of their assigned duties (e. g., Metcalf et al., 2019), and many who have not. Interview questions were tailored to context and did not adopt consistent predetermined structure(Seidman, 2006): some portion of participants spoke to me about their own ethical concerns, often after raising them to their employers and seeing them go unaddressed. This study received IRB or other ethics approval as appropriate.
@@
The variety of contexts in which I investigate AI ethics practices allows us to understand how a variety of epistemological stances interact with the legitimacy of AI ethics claims, and what is rendered possible or not as a result of these stances. Participants were invited to self-identify their gender: 57 identified as men, 13 identified as women, four identified as non-binary, and one declined to answer. Most interviews lasted one hour (though ranging between 25 minutes and 3 hours), and were transcribed and analyzed using inductive thematic analysis (Braun and Clarke, 2006, 2019) [^3] In an inductive thematic analysis, themes are grounded in and emerge from the data, as opposed to deductive analysis where themes are decided a priori, and is properly a reflexive process, where the researcher reflects on their assumptions, training, values, and how this may shape analysis. See also: (Braun and Clarke, 2021, 2006)., under an interpretivist epistemological paradigm (Lincoln and Guba, 2000).444 Interpretivism, similar to its younger cousin constructivism (Schwandt, 1994), understands that individuals are experts in interpreting their own experience, asks researchers to deeply understand these interpretations, and with both sharing the task of constructing wider meaning from these. To positivists, this is an unacceptable limitation, as no view from nowhere objectivity is possible. See also (Lincoln and Guba, 2000; Daston and Galison, 2007; Gebru, 2021; Haraway, 1988). My analysis is therefore inflected by my own positionality: I am a male-identified scholar in the United States with formal training in computer science, mentored by feminist STS scholars, with experience organizing against pervasive surveillance (Guo and Ryan-Mosley, 2023; Gulland and Tan, 2023) and carceral technology (Davidson, 2020; Kubit, 2022). It is from attempts to reconcile differences in language and legitimacy between the first and last of these–computer science and carceral tech activism–that I am motivated to write this paper, but my positionality necessarily limits those I can talk with and inflects the interpretations I arrive at, and others may arrive differently. In some ways, these locations represent the “stepping in and stepping out” characteristic of fieldwork (Powdermaker, 1966) in that some of my nonbinary and woman-identified participants construct their AI ethics practice as self-consciously feminist or even necessarily constructed in femme-only spaces, yet also, the predominantly technical and male backdrop in which AI ethics work sits paints me as part of the ingroup relative to the bulk of participants in this work: men in technical roles.
@@
2. The lower status of ethics work within AI cultures
This section demonstrates how my participants understand ethics work to be lower status relative to “real” AI work, but similar phenomena have been examined before. Past research shows how certain work in technology communities, like writing documentation or building community, is often done by women, and is less valued than writing code (Nafus, 2012). More recent work within AI has even shown differences between high status work on models versus lower status work on the dataset models are built from (Sambasivan et al., 2021), and recent manifestos have attacked “tech ethics” on the basis that it hinders “progress” on endeavors such as “Becoming Technological Supermen” (Andreessen, 2023). Forsythe shows how these status divides are inscribed in the language used by the AI practitioners she studied: they restricted their use of the word “‘work’ to what they think of as ‘real AI’: building systems”  (Forsythe, 2001, p. 26), set in contrast to the other kinds of work such as grant writing, meetings, and also the “knowledge elicitation” interviews seen as boring but necessary to “capture” knowledge to encode into systems.
@@ 6. Conclusions: towards humble technical practices
Epistemic power is bound up with, and often flowing from, other kinds including economic and political power. For example, Forsythe writes how the (poor) appropriation of ethnographic methods in positivist ways by untrained AI researchers had economic outcomes: “my own research funds have dried up” (Forsythe, 2001, p. 143). It is therefore those with other kinds of power, along lines of race, gender, or wealth, that should first share or cede epistemic power. Notably, people harmed by AI are often not paid nor recognized for sharing their lived experience of this harm in the ways that academics may be paid to collect it (Sloane et al., 2022). Therefore, I call for more ways to legitimize the sharing of lived experience in and on its own terms as valid AI ethics practice, and indeed, the preeminent and most true form of such practice. This must be listened to, and AI Ethics must be seen as a project to legitimize and amplify these on their own terms.